### v1

# Misquoted No More: Securely Extracting F Programs with IO

Abstract.
Shallow embeddings that use monads to represent effects are popular in proof-oriented languages because they are convenient for formal verification. Once shallowly embedded programs are verified, they are often extracted to mainstream languages like OCaml or C and linked into larger codebases. The extraction process is not fully verified because it often involves quotation—turning the shallowly embedded program into a deeply embedded one—and verifying quotation remains a major open challenge. Instead, some prior work obtains formal correctness guarantees using translation validation to certify individual extraction results. We build on this idea, but limit the use of translation validation to a first extraction step that we call relational quotation and that uses a metaprogram to construct a typing derivation for the given shallowly embedded program. This metaprogram is simple, since the typing derivation follows the structure of the original program. Once we validate, syntactically, that the typing derivation is valid for the original program, we pass it to a verified syntax-generation function that produces code guaranteed to be semantically related to the original program.
We apply this general idea to build SEIO, a framework for extracting shallowly embedded F programs with IO to a deeply embedded $\lambda$-calculus while providing formal secure compilation guarantees. Using two cross-language logical relations, we devise a machine-checked proof in F that SEIO guarantees Robust Relational Hyperproperty Preservation (RrHP), a very strong secure compilation criterion that implies full abstraction as well as preservation of trace properties and hyperproperties against arbitrary adversarial contexts. This goes beyond the state of the art in verified and certifying extraction, which so far has focused on correctness rather than security.
secure compilation, formal verification, proof assistants, input-output, shallow embeddings, monads, metaprogramming, quotation, extraction, translation validation, logical relations
copyright: ccccs: Software and its engineering Software verificationccs: Software and its engineering Compilers
1. Introduction
The F proof-oriented language (Swamy et al., 2016) is commonly employed as a programming language, wherein programs are shallowly embedded using monads to represent effects (Ahman et al., 2018; Maillard et al., 2019; Ahman et al., 2017; Rastogi et al., 2021). Working with a shallow embedding eases formal verification, after which such programs are extracted to a mainstream language like OCaml (Letouzey, 2004) or C (Protzenko et al., 2017). One success story of this approach is EverCrypt (Protzenko et al., 2020): a cryptographic provider written and verified in F and then extracted to C code that was integrated in Mozilla Firefox, the Linux kernel, and WireGuard (Ahman et al., 2025).
Extraction of F programs is, however, unverified and part of the trusted computing base. Moreover, in case the extracted code gets linked into a bigger unverified codebase, then the unverified code is also trusted. This is problematic because the unverified code could inadvertently or maliciously (if compromised by an attacker) break the internal invariants of the verified code, leading to unsoundness and insecurity. In other words, verified F programs lose their formal guarantees when extracted and linked with unverified code.
To address this, we aim to provide formal security guarantees for extracted F code even when linked with arbitrary unverified code. We are not aware of any proof-oriented language whose extraction provides such strong secure compilation guarantees (Patrignani et al., 2019; Abate et al., 2019). Most existing work focuses on verifying correctness of extraction (Forster et al., 2024; Meier et al., 2025; Anand et al., 2017; Annenkov et al., 2022; Nezamabadi et al., 2026; Hupel and Nipkow, 2018), not security, and, moreover, extraction is not fully verified. To fully verify extraction of shallowly-embedded programs, one would have to define the programs, define extraction, and connect the two, all within the same language. This represents a significant challenge: the programs are shallowly embedded, whereas the extraction function operates on a deep embedding (Forster et al., 2024). Connecting the two requires a metaprogram, and the most direct way of obtaining formal guarantees would require verifying this metaprogram. This is challenging, for a start because verifying a metaprogram requires a mechanized meta-theory for the proof-oriented language, but there has been steady progress towards this goal for many such languages (Carneiro, 2024; Liesnikov and Cockx, 2024; Hupel and Nipkow, 2018; Nezamabadi et al., 2026; Cohen and Johnson-Freyd, 2024; Sozeau et al., 2025). Another challenge though is that the standard primitive connecting a shallow and a deep embedding is quotation—a metaprogram that translates a shallowly embedded program into its deep embedding. As far as we know, quotation was not verified for any proof-oriented language, and even for Rocq, which has a mature mechanized meta-theory (Sozeau et al., 2025), verifying quotation remains a big unsolved challenge. [^1] Verification of quotation was started in MetaRocq (Sozeau et al., 2025), but (as of 19 February 2026) it has not progressed much in the last 3 years: https://github.com/MetaRocq/metarocq/tree/9.1/quotation/theories
Instead of verifying extraction, other works certify each extraction result using translation validation to obtain formal correctness guarantees (Pit-Claudel et al., 2022; Myreen and Owens, 2014; Abrahamsson et al., 2020; Mullen et al., 2018; Forster and Kunze, 2019). A metaprogram is used on a shallowly embedded program to generate both the extracted code and a proof that the two are related. This is an instance of translation validation because to ensure trust each such proof has to be checked by the proof-oriented language (e.g., type-checked), which can also fail because the metaprogram is not verified and the proof can be wrong too or fail to verify (e.g., fail during unification or SMT solving), thus, ideally, one would limit the use of translation validation.
In this paper, we introduce SEIO, a formally secure extraction framework for F, written itself in F. SEIO advances the state of the art on certifying extraction (Pit-Claudel et al., 2022; Myreen and Owens, 2014; Abrahamsson et al., 2020; Mullen et al., 2018; Forster and Kunze, 2019) in two key ways: (1) it minimizes the reliance on translation validation, and (2) it provides formal secure compilation guarantees against arbitrary adversarial contexts. Specifically, SEIO extracts IO—a shallow embedding of simply typed programs capable of file-based Input-Output (IO)—to $\lambda_{io}$, a deep embedding of a terminating call-by-value $\lambda$-calculus with IO primitives (Figure 1).
 $\small\begin{array}[]{cc}\textit{\ verified program}\end{array}$$\small\begin{array}[]{cc}\textit{typing derivation}\end{array}$extractedprogramunverifiedcodetarget language ($\lambda_{io}$)source language (IO)syntaxgenerationlink relationalquotation
Figure 1. An overview of extraction using SEIO. Solid lines are verified pure F functions. The dashed arrow represents the relational quotation metaprogram, the result of which is checked by F to be a typing derivation for the original verified program.
Contributions:
[…]
    We introduce relational quotation, a translation validation technique for quoting shallowly embedded programs in a dependently typed language. The idea is to define a typing relation that precisely characterizes the shallowly embedded source language, and to use a metaprogram to construct a typing derivation for a given program by following its structure. Checking that a derivation is indeed a valid typing derivation for the original program actually happens by type checking the derivation. This works because the typing relation uses dependent types to guarantee that a derivation corresponds to a specific shallowly embedded program. From a valid derivation, we can always extract the program correctly, since the rest of extraction can be defined as a pure function for syntax generation that is verified once and for all. Because the type of the derivation ties it to the original program, we can semantically relate the generated code back to that program.
[…]
    We demonstrate this general idea by extracting IO programs with SEIO. For this, we show how to relate a shallowly embedded IO program with a deeply embedded $\lambda_{io}$ one using two cross-language logical relations (Neis et al., 2015; Ahmed, 2015): one that semantically relates the extracted program to the original shallowly embedded source program, and one that does the opposite.
[…]
    We use the logical relations to provide a machine-checked proof that SEIO satisfies Robust Relational Hyperproperty Preservation (RrHP). This is the strongest secure compilation criterion of Abate et al. (2019), which not only implies full abstraction (i.e., preserving observational equivalence), but also the preservation of trace properties and hyperproperties (like noninterference) against adversarial contexts. Since RrHP is transitive, we expect that in the future SEIO can be vertically composed with previous secure compilation steps for F that were proved to satisfy RrHP, but only down to some shallowly embedded languages (Andrici et al., 2024, 2025) (§ 8).
Outline. § 2 starts by illustrating the key ideas of SEIO on a simple running example. § 3 details how relational quotation works for IO. § 4 formalizes the IO and $\lambda_{io}$ languages, and the two logical relations between them. § 5 shows how all the pieces of SEIO come together and how we use the logical relations to prove that it satisfies RrHP. § 6 provides more details on our running example, and discusses how we further compile $\lambda_{io}$ in practice using the Peregrine framework (Chapman et al., ). Finally, we discuss related work in § 7, and conclude and mention future work in § 8.
Artifact. This paper comes with an artifact in F that contains SEIO and a complete machine-checked proof that it satisfies RrHP. The proofs of some of the low-level lemmas we stated in this formalization were completed by Claude Opus 4.6, which is extremely good at doing F proofs with minimal human guidance (Swamy, 2026). The artifact includes over 900 top-level definitions (including type and function definitions, lemmas, and test programs) and for around 75 of them the actual definition was produced by Claude, usually lemmas for which we had written the statement and Claude produced the complete proof including by adding additional lemmas in a vibe proving session that took a couple of days and would cost a Claude user less than 15\$. The largest such proofs are over 200 lines long, look natural to us, and would have taken us much longer to write. The artifact is available at: https://github.com/andricicezar/fstar-io/tree/icfp26/seiostar
2. Key Ideas
We present the main ideas of SEIO using the simple running example of a verified wrapper for unverified AI agents. We first give an overview of SEIO (§ 2.1), then we present relational quotation (§ 2.2), and finally we present what guarantees SEIO provides when extracting the wrapper (§ 2.3).
1let validate : string -> task_t -> string -> bool =
2
3let wrapper (f: string) (task: task_t) (agent: string->task_t->io unit) : io (either unit err) =
4 let! contents = read_file f in
5 agent f task;!
6 let! new_contents = read_file f in
7 if validate contents task new_contents then io_return (Inl ()) else io_return (Inr Failed_validation)
Figure 2. Running example: a validation wrapper for unverified agents.
2.1. Overview of SEIO
In Figure 2 (lines 3-7) we implement a validation wrapper for unverified AI agents. [^2] F syntax (Andrici et al., 2024) is similar to OCaml (val, let, match, etc). Binding occurrences b take the form x:t or #x:t for an implicit argument. Lambda abstractions are written fun b_1 bn -> t, whereas b_1 -> -> bn -> C denotes a curried function type with result C. Contiguous binders of the same type may be written (v_1 vn : t). Dependent pairs are written as x:t_1 & t2. For non-dependent function types, we omit the name in the argument binding, e.g., type #a:Type -> (#m #n : nat) -> vec a m -> vec a n -> vec a (m+n) represents the type of the append function on vectors, where both unnamed explicit arguments and the return type depend on the implicit arguments. Type0 is the lowest universe of F; we also use it to write propositions, including True (True) and False (False). We generally omit universe annotations. The type either t_1\$~\$t_2 has two constructors, Inl:(#t_1 #t_2:Type) -> t_1 -> either t_1\$~\$t_2 and Inr:(#t_1 #t_2:Type) -> t_2 -> either t_1\$~\$t_2. Expression Inr? x tests whether x is of the shape Inr y. Binary functions can be made infix by using backticks: x op\$~\$y stands for op x y. Finally, when writing explicitly monadic code one writes let! x = e_1 in e_2 for bind. Such a wrapper is useful for tasks where it is easier to check the validity of the result than doing the task itself. An agent is supposed to perform such a task on a file f, and the wrapper validates that the agent performed the task correctly. The wrapper saves the initial contents of the file (line 4), then calls the agent (line 5), and finally checks that the task was performed correctly (lines 6-7) using a function validate. If calling the wrapper is successful (denoted by returning Inl ()), then the new contents of the file correspond to correctly performing the task on the original file.
The wrapper is written in the io monad, which includes operations such as openfile, read, write, close, etc. Each of these operations can fail. The function read_file used in the Figure 2 just opens a file, reads its contents, and closes it back. At this level the agent is modeled as a computation in the io monad, but it will actually be implemented in $\lambda_{io}$.
The wrapper is shallowly embedded in F and we want to define extraction such that it produces a syntactic expression in $\lambda_{io}$, which is a deep embedding of a call-by-value (CBV) $\lambda$-calculus with IO primitives.
let extracted_wrapper =
ELambda (ELambda (ELambda
eLet (eReadFile (EVar 2)) (
eLet (eApp 2 (EVar 1) (EVar 3) (EVar 2)) (
eLet (eReadFile (EVar 4)) (
EIf (eApp 3 eValidate (EVar 2) (EVar 5) (EVar 0)) (EInl Ett) (EInr EFailedValidation))))))
The eLet, eReadFile, eApp x and eValidate are combinators that help with readability.
One can generate extracted_wrapper starting from wrapper only by using a metaprogram. Existing work on certifying extraction (Pit-Claudel et al., 2022; Myreen and Owens, 2014; Abrahamsson et al., 2020; Mullen et al., 2018; Forster and Kunze, 2019) would use a metaprogram to generate both extracted_wrapper and a proof that the wrapper and the extracted_wrapper are semantically related, proof that has to be type-checked.
The main insight that distinguishes SEIO (Figure 1) from previous work is that the metaprogram does not have to directly produce a deeply embedded program. Instead, SEIO uses a metaprogram to produce a typing derivation for the original shallowly embedded IO program, which we call relational quotation. Once we validate, syntactically, that the typing derivation is valid for the original program extraction continues with a pure F function that generates $\lambda_{io}$ syntax, which we verify once and for all to produces code guaranteed to be semantically related to the original program. This makes SEIO rely less on translation validation and more on verification compared with existing work (Pit-Claudel et al., 2022; Myreen and Owens, 2014; Abrahamsson et al., 2020; Mullen et al., 2018; Forster and Kunze, 2019), and is further detailed in the next subsection.
2.2. Relational quotation
Relational quotation is a technique to implement quotation for a subset of a dependently typed language that uses translation validation to ensure correctness. Relational quotation involves using dependent types to define a typing relation that carves out the subset of the proof-oriented language that represents the shallowly embedded source language. For example, for a simply typed pure language shallowly embedded in F the typing relation would have the following signature:
type typing_oval : tyG:typ_env -> a:Type -> (eval_env tyG -> a) -> Type
As we are dealing with a shallow embedding, this typing relation is not defined over syntactic types and expressions. Instead, the typing relation relates a typing environment (a partial map from variables to F types, denoted by tyG), an F type (a), and an open F value, which is a function from an evaluation environment (a map from variables to F values) to a value of type a.
To understand why we need to work with open F values, consider we want to build a typing derivation for the identity function on ints fun (x:int) -> x. We need a way to refer to the body of the function, and usually one would just have the variable x as the body, but this is not possible when working with a shallow embedding because we cannot have a free variable x as an F value. This is why we “open” F values by assuming that we have an evaluation environment fsG that contains the free variables. We do this by representing open values as functions from evaluation environments to values (eval_env tyG -> a), where variables are de Bruijn indices, the evaluation environment is guaranteed to contain a value for each entry in the typing environment (because of the type eval_env tyG), and the evaluation environment behaves like a stack, with the usual push, hd and tail operations. For example, for a typing environment tyG = extend int empty, we can represent the body of the identity function as body = fun fsG -> hd fsG. Further, we represent the identity function as fun fsG x -> body (push fsG x), which, by unfolding the body, simplifies to fun fsG x -> hd (push fsG x), which reduces to fun _ x -> x, which is syntactically equal to the (thunked) identity function. Open F values enable us to define typing rules for free variables, which is key to define a typing relation that supports lambda abstractions and characterizes a shallowly embedded programming language.
Our typing rules resemble standard typing rules, but instead of syntax expressions, they operate on F types and open F values. For example, the rule for introducing the literal false (Qfalse), types F’s false as a bool, where bool is an F type.
type typing_oval : tyG:typ_env -> a:Type -> (eval_env tyG -> a) -> Type =
| Qfalse : #tyG:typ_env ->
typing_oval tyG bool (fun _ -> false) (* this is F*’s bool and false *)
| QVar0 : #tyG:typ_env -> #a:Type ->
typing_oval (extend a tyG) a (fun fsG -> hd fsG)
| QVarS : #tyG:typ_env -> #a:Type -> #b:Type -> #x:(eval_env tyG -> a) ->
typing_oval tyG a x ->
typing_oval (extend b tyG) a (fun fsG -> x (tail fsG))
| QApp : #tyG:typ_env -> #a:Type -> #b:Type -> #f:(eval_env tyG -> a -> b) -> #x:(eval_env tyG -> a) ->
typing_oval tyG (a -> b) f ->
typing_oval tyG a x ->
typing_oval tyG b (fun fsG -> (f fsG) (x fsG))
| QLambda : #tyG:typ_env -> #a:Type -> #b:Type -> #body:(eval_env (extend a tyG) -> b) ->
typing_oval (extend a tyG) b body ->
typing_oval tyG (a -> b) (fun fsG x -> body (push fsG x))
| QInl : #tyG:typ_env -> #a:Type -> #b:Type -> #v:(eval_env tyG -> a) ->
typing_oval tyG v ->
typing_oval tyG (either a b) (fun fsG -> Inl (v fsG))
(* Qtrue, QIf, Qtt, QInr, QCase, etc., are similarly defined *)
The rule QVar0 introduces the most recently bound variable (index 0) by projecting the head of the evaluation environment. QVarS weakens the environment by discarding the most recent binding via tail and recursing on the shorter environment. The QApp rule types F’s function application by requiring that f and x are well-typed in the same environment and producing the open value fun fsG -> (f fsG) (x fsG), which applies the function to its argument pointwise over the environment. The QLambda rule requires typing the body in an environment extended with a fresh binding of type a. The resulting open value fun fsG x -> body (push fsG x) pushes the argument onto the evaluation environment before evaluating the body.
For the simplicity of exposition, the typing rules above were given for simply typed pure F code. To also support typing of code that uses the io monad, we actually follow the typing rules of the fine-grained call-by-value $\lambda$-calculus (FGCBV) (Levy et al., 2003), which makes a clear distinction between values and potentially effectful computations. Intuitively, values are terms in introduction forms (which do not evaluate further by themselves, such as $\lambda$–abstractions, Boolean values, and pairs of values), and computations contain different kinds of elimination forms (such as function application, conditional statements, and pattern matching), the monadic operations of returning values and sequentially composing multiple computations, and file operations. The typing relations for values and computations are then defined by mutual induction, and have the following signatures:
type typing_oval : tyG:typ_env -> a:Type -> (eval_env tyG -> a) -> Type
type typing_oprod : tyG:typ_env -> a:Type -> (eval_env tyG -> io a) -> Type
We present the full definition of these typing relations in Figure 3 in § 3.1. We call the language carved out by these typing relations IO. Next, we present a typing derivation for the wrapper that proves that it is written in IO.333qLet, qReadFile, qValidate, qApp x and qVar x are helper functions to simplify readability.
As wrapper is a closed program, the typing judgement is applied to the empty typing environment.
let wrapper_typing : typing_oval empty _ (fun _ -> wrapper) =
QLambda (QLambda (QLambdaIO
qLetIO (qReadFile (qVar 2)) (
qLetIO (QAppIO (QApp (qVar 1) (qVar 3)) (qVar 2)) (
qLetIO (qReadFile (qVar 4))
QIfIO (qApp 3 qValidate (qVar 2) (qVar 5) QVar0)
(QReturn (QInl Qtt))
(QReturn (QInr QFailedValidation))))))
There are three important things to notice about this derivation. First, the type of the derivation is typing_oval empty _ (fun _ -> wrapper), which directly relates the derivation with the original wrapper. Second, we do not have to spell out the implicit arguments for typing environments, types and open F values/computations (remember that in F definitions such as typing such implicit arguments are preceded by #). Third, the structure of the derivation precisely follows the structure of wrapper.
To type check such a derivation, we rely on F inferring all the implicit arguments, which given the structure of our the typing relation F should always be able to do. The intermediate typing environment and F values in the derivation are inferred with help from the top-level type we specify for the whole derivation (e.g., if QInl Qfalse is typed as typing_oval empty (either bool unit) (fun _ -> Inl false), then it gets elaborated to QInl #empty #bool #unit (Qfalse #empty)). Since we work with simple types, it should always be possible to infer these implicits. The more complex implicits for open F values can be inferred because of the structure of the typing rules that propagates them bottom-up. The bottom-up propagation of open values constructs a program that at the end has to be equal to the wrapper, and only if that is the case does the derivation type-check. F can unify the two programs because the derivation follows the syntactic structure of the program, and because the passing of the evaluation environment around gets simplified away. We already saw how the simplification happens when we looked at the identity example: during unification, expression such as tail (push x fsG) and hd (push x fsG) are simplified to fsG and x. The evaluation environment can always be simplified away when one tries to type a (closed) F value because the empty environment is used.
Relational quotation is a form of translation validation because the derivation has to be type-checked and we cannot formally guarantee that the derivations generated by the metaprogram will type-check (without verifying both the metaprogram, which includes the standard F quotation, and the F type-checker itself). This doesn’t increase the TCB, since the F type-checker is already trusted for verifying the original shallowly embedded programs.
The metaprogram necessary to produce a typing derivation is rather simple (§ 3.2). It takes an F program (e.g., wrapper), quotes it using standard F quotation, and produces a typing derivation (e.g., wrapper_typing) by following the structure of the program any applying the corresponding typing rules. The fact that we did not had to spell out the implicits for the manual derivation of the wrapper above is also true for the metaprogram, since the implicits get instantiated automatically when the type-checking of the derivation. Just by type-checking, the derivation is checked to be a valid typing derivation for the original program, which is a syntactic match.
It is great that relational quotation needs a simple metaprogram because this metaprogram is not verified. Previous work (Pit-Claudel et al., 2022; Myreen and Owens, 2014; Abrahamsson et al., 2020; Mullen et al., 2018; Forster and Kunze, 2019) on certifying extraction produces both a deep embedding and a proof. Moreover, at first sight, wrapper_typing looks similar to extracted_wrapper, but we deal with two different languages: IO is a FGCBV language, while $\lambda_{io}$ is a CBV language. Previous work deals with the complexity of semantically relating two different language semantics during translation validation, while in SEIO this only happens in the verified syntax generation step (§ 5).
2.3. Formally secure extraction from IO to $\lambda_{io}$
From the typing derivation, it is very easy to produce extracted_wrapper. SEIO has a pure F function that goes recursively over the derivation and produces the expected $\lambda_{io}$ program. Each constructor of the typing relation corresponds directly to a syntactic form in $\lambda_{io}$: e.g., QBind maps to a let-binding, QOpenfile to a call to the openfile primitive, and QLambda / QLambdaIO to lambda abstractions.
To relate IO programs with $\lambda_{io}$ programs, we give trace-producing semantics to IO and $\lambda_{io}$, meaning that for each IO operation, we associate an event. The events include both the argument and the result of an operation. The events get appended in order to build a local trace. A possible successful local trace for the wrapper is the following, where fd and fd are universally quantified, but guaranteed to be fresh.
[EvOpenfile f (Inl fd); EvRead fd (Inl contents); EvClose fd; (** first read_file **)
agents trace …;
EvOpenfile f (Inl fd’); EvRead fd (Inl new_contents); EvClose fd’] (** second read_file **)
We can use the trace-producing semantics of IO to verify the wrapper, since it allows us to write postconditions over a final result and a local trace of events. We can verify the wrapper to satisfy a safety property in arbitrary contexts: which states that for any file, task and agent, if the wrapper runs successfully (the Inl case), looking at the trace of events, the contents last read from the file corresponds (according to the validate function) to performing the task on the contents first read from the file.
[…]
Since SEIO guarantees that whenever extraction succeeds it preserves safety against adversarial contexts we can get that the extracted_wrapper also satisfies this safety property when it is linked with arbitrary unverified agents.
type typing_oval : tyG:typ_env -> a:Type -> (eval_env tyG -> a) -> Type =
| QCase : #tyG:typ_env -> #a:Type -> #b:Type -> #c:Type ->
#sc:(eval_env tyG->either a b) -> #il:(eval_env (extend a tyG)->c) -> #ir:(eval_env (extend b tyG)->c) ->
typing_oval tyG (either a b) sc -> typing_oval (extend a tyG) c il -> typing_oval (extend b tyG) c ir ->
typing_oval tyG c (fun fsG -> match sc fsG with | Inl x -> il (push fsG x) | Inr y -> ir (push fsG y))
| QLambdaIO : #tyG:typ_env -> #a:Type -> #b:Type -> #body:(eval_env (extend a tyG) -> io b) ->
typing_oprod (extend a tyG) b body ->
typing_oval tyG (a -> io b) (fun fsG x -> body (push fsG x))
(* Qfalse, QVar0, QVarS, QApp, QLambda were presented in 2.2 *)
(* Qtt, Qtrue, QStringLit, QIf, QMkpair, QFst, QSnd, QInl, QInr are similarly defined *)
and typing_oprod : tyG:typ_env -> a:Type -> (eval_env tyG -> io a) -> Type =
| QOpenfile : #tyG:typ_env -> #fnm:(eval_env tyG -> bool) ->
typing_oval tyG string fnm ->
typing_oprod tyG (either file_descr err) (fun _ -> io_openfile fnm)
(* QRead, QWrite, QClose are similar *)
| QReturn : #tyG:typ_env -> a:Type -> #x:(eval_env tyG -> a) ->
typing_oval tyG a x ->
typing_oprod tyG a (fun _ -> io_return x)
| QBind : #tyG:typ_env->#a:Type->#b:Type -> #m:(eval_env tyG->io a) -> #k:(eval_env (extend a tyG)->io b) ->
typing_oprod tyG a m -> typing_oprod (extend a tyG) b k ->
typing_oprod tyG b (fun fsG -> io_bind (m fsG) (fun x -> k (push fsG x)))
| QAppIO : #tyG:typ_env -> #a:Type -> #b:Type -> #f:(eval_env tyG -> a -> io b) -> #x:(eval_env tyG -> a) ->
typing_oval tyG (a -> io b) f -> typing_oval tyG a x ->
typing_oprod tyG b (fun fsG -> (f fsG) (x fsG))
| QCaseIO : #tyG:typ_env -> #a:Type -> #b:Type -> #c:Type ->
#sc:(eval_env tyG->either a b) -> #il:(eval_env (extend a tyG)->io c) -> #ic:(eval_env (extend b tyG)->io c) ->
typing_oval tyG (either a b) sc -> typing_oprod (extend a tyG) c il -> typing_oprod (extend b tyG) c ic ->
typing_oprod tyG c (fun fsG -> match sc fsG with | Inl x -> il (push fsG x) | Inr y -> ic (push fsG y))
(* QIfIO is similar to QCaseIO *)
Figure 3. The typing relation for IO.
3. Relational quotation of IO
This section details how relational quotation works for IO. We first present the typing relation that carves out the IO subset of F (§ 3.1), and then describe the metaprogram that generates typing derivations (§ 3.2).
3.1. Typing relation
The typing relation operates over a fixed set of simple F types. These include base types (unit, bool, string, and file_descr), pure function types (a -> b), effectful function types (a -> io b), product types (a & b), and sum types (either a b). The full typing relation is given in Figure 3. Following FGCBV (Levy et al., 2003), the relation is split into two mutually defined judgements: typing_oval for values and typing_oprod for io computations.
The typing relation is defined on the interface of the io monad, which provides io_return and io_bind with their standard types, and a few primitives for file operations: io_openfile (open a file by name), io_read / io_write (read from / write to a file descriptor), and io_close (close a file descriptor). Each operation can fail, so e.g., io_openfile returns io (either file_descr err).
Values (typing_oval).
The value typing rules from § 2.2 (Qfalse, QVar0, QVarS, QApp, QLambda) are complemented with analogous rules for other introduction forms: constants (Qtrue, Qtt, QStringLit), constructors (QInl, QInr, QMkpair), and projections (QFst, QSnd). Two rules deserve special attention. First, QCase allows pattern matching on either a b when both branches produce a value: the scrutinee is typed as a value, and each branch is typed in the environment extended with the matched component. QIf is similar for Booleans. In the FGCBV terminology of Levy et al. (2003), these are complex values: elimination forms whose result is a value rather than a computation. Second, QLambdaIO is the rule for effectful $\lambda$-abstractions: it types a value of type a -> io b by requiring the body to be a well-typed io computation.
Computations (typing_oprod).
The rules for typing io computations include: QReturn for monadic return, QBind for sequencing (where the continuation is typed in the environment extended with the intermediate result), and IO operations (QOpenfile, QRead, QWrite, QClose), each of which requires its argument to be a well-typed value. The remaining rules handle elimination forms whose result is a computation: QAppIO for applying an effectful function (a -> io b), and QCaseIO (resp. QIfIO) for pattern matching (resp. conditionals) where the branches are computations. Note that case analysis and conditionals appear in both relations—QCase/QIf for values and QCaseIO/QIfIO for computations—because the same elimination can yield either kind of term depending on the context.
Expressivity of IO.
The language carved out by these typing relations is a simply typed FGCBV language with IO primitives. It supports higher-order functions (both pure and effectful), pairs, sums, Booleans, strings, and file descriptors. All IO programs are terminating. Supporting refinement types, dependent types, recursive types, or recursion is left as future work (§ 8).
3.2. Generating derivations using a metaprogram
We write our relational quotation metaprogram to produce derivations directly in F, which supports metaprogramming by providing a primitive effect and a set of primitives that allow metaprograms to interact with F’s type-checker via an API called Meta-F (Martínez et al., 2019). In this section, we present how the key ideas about the metaprogram (§ 2.2) are realized.
We define a metaprogram generate_derivation that takes the quotation of an F program, and tries to create a derivation for it. If the program is not in the IO language, then the metaprogram will fail. To produce the quotation of the wrapper (Figure 2), we use the backtick operator, which is a special F quotation primitive for top-level definitions (wrapper). To call the instantiated metaprogram inside F, we use the special splice_t primitive, which expects the metaprogram to return a list of F terms representing new top-level let bindings. It requires that the new bindings are well typed; and in this case there is just one with new binding named wrapper_typing.
%splice_t[wrapper_typing] (generate_derivation (‘wrapper))
The generate_derivation metaprogram does the following steps:
  1. (1)
    Goal Formation: It constructs the expected type for the derivation by using the empty typing environment and thunking the program: typing_oval empty _ (fun _ -> wrapper).
  2. (2)
    Term Generation: It generates a term for the derivation with uninstantiated implicit arguments. This is a syntactic mapping from F syntax to the typing relation constructors. The mapping distinguishes between pure terms and computations in the io monad, selecting appropriate typing rules. We have to keep track of the types of the binders so that we know when to use QApp or QAppIO. We map F’s locally nameless representation to our representation, and we inline top-level definitions.
  3. (3)
    Type-checking: Finally, we have to prove that the derivation term has the expected type (from step 1). Using Meta-F, we split it into three steps to have more control over proving the type equality:
    1. (a)
      Implicit Argument Instantiation: It calls Meta-F to infer the implicit arguments of the generated term (from step 2) using the expected top-level type (from step 1). This is one call and no special guidance is necessary for inference to be complete. The unification of the typing environment and of the types should be decidable, since the typing rules correspond to a simply typed FGCBV language. The unification of open values and computations should always succeed because of the structure of the typing rules, which propagate them bottom-up.
    2. (b)
      Typing inference: At this moment, Meta-F does not know yet the type of the derivation term. It got instruction just to instantiate the implicit arguments, which could be incomplete. Since we know the instantiation should be complete, the metaprogram calls Meta-F to infer the type of the derivation term.
    3. (c)
      Type Equality Check: It checks that the inferred type of the derivation matches the expected type (from step 1). Crucially, this step checks that the program constructed by the derivation is equal to the original program (wrapper). To solve the equality, it manually unfolds some specific definitions, and then it uses F’s reflexivity tactic (trefl), which first applies F’s normalization to unfold definitions and to reduce primitive operations in both programs, and then performs unification. The equality check should always succeed because the derivation follows the structure of the program.
4. Relating the trace-producing semantics of IO and $\lambda_{io}$
In this section we formally relate the shallowly embedded IO language with the deeply embedded $\lambda_{io}$ language that IO programs are extracted to. $\lambda_{io}$ is an ML-like language based on the simply typed $\lambda$-calculus. We relate it to IO by equipping both languages with trace-producing semantics.
4.1. Syntax of $\lambda_{io}$
We begin by defining the syntax of $\lambda_{io}$ expressions as the inductive F type exp.
type exp =
[…]
| EFileDescr : file_descr -> exp (* file descriptor *) | EString : string -> exp (* string constant *)
| ERead : exp -> exp (* reading from a file *) | EWrite : exp -> exp -> exp (* writing to a file *)
| EOpen : exp -> exp (* opening a file *) | EClose : exp -> exp (* closing a file *)
(* there are also constructors for EUnit, Etrue, Efalse, EApp, EIf, EPair, EFst, ESnd, ECase, EInl, EInr *)
We also recursively define a predicate is_value e that identifies expressions e which are values, i.e., that are closed expressions in introduction form, such as $\lambda$-abstractions and pairs of values.
The ERead, EWrite, EOpen, and EClose expressions represent the extension of the simply typed $\lambda$-calculus to include effectful IO operations. These operations can fail. Their results are of type either a err, where the left injection represents successful execution and the right injection of error represents a failed execution. We use strings to represent file names and the data that can be read from or written to a file. We use file descriptors to represent the unique descriptor generated by the operating system with each open file call. For simplicity, these file descriptors are modelled as natural numbers. ERead takes a file descriptor and returns an optional string, EWrite takes a file descriptor and a string and returns an optional unit, EOpen takes a string and returns an optional file descriptor, and EClose takes a file descriptor and returns an optional unit.
4.2. Small-step trace-producing operational semantics of $\lambda_{io}$
Next, we equip $\lambda_{io}$ with an operational semantics. Our semantics is based on sequences of events, called traces, which arise from the execution of IO operations (ERead, EWrite, EOpen, and EClose). Events are indexed by the arguments and result of the operation, e.g., executing the read operation ERead (EFileDescr fd) produces an event EvRead fd r, where r is the optional string that was read.
To link the computations being executed to the traces they produce, we represent the small-step operational semantics of $\lambda_{io}$ as a reduction relation defined as an inductive type indexed by two $\lambda_{io}$ expressions (representing one step of execution), a list of past events (history), and an optional event indexed by the history. The optional event allows us to explicitly differentiate between effectful and pure reduction steps. Pure reduction steps are indexed by None, and reduction steps involving the execution of IO operations are indexed by the corresponding events. As is standard practice, we define the computation rules of the operational semantics on closed expressions (Ahmed, 2004). The reduction rules include congruence rules to transform subexpressions into value forms, and computation rules to execute expressions when they have been reduced to redex forms. To illustrate the step relation, we present one pure reduction rule ($\beta$-reduction), and the congruence rule and two reduction rules for the (un)successful execution of file opening. Reduction rules for other expressions (in particular, other IO operations) are defined analogously, and can be found in the artifact.
type step : closed_exp -> closed_exp -> h:history -> option (event_h h) -> Type =
| SBeta : e11:exp{is_closed (ELam e11)} -> e2:value -> h:history ->
step (EApp (ELam e11) e2) (subst_beta e2 e11) h None
| SOpen : #str:closed_exp -> #str’:closed_exp -> #h:history -> #oev:option (event_h h) ->
step str str h oev ->
step (EOpen str) (EOpen str’) h oev
| SOpenReturnSuccess : str:string -> h:history ->
step (EOpen (EString str)) (EInl (EFileDescr (fresh_fd h))) h (Some (EvOpen str (Inl (fresh_fd h))))
| SOpenReturnFail : str:string -> h:history ->
step (EOpen (EString str)) (EInr (EString err”)) h (Some (EvOpen str (Inr err”)))
We note that the optional event argument of step is indexed by the history because certain effectful steps use the history to generate the $\lambda_{io}$ expression to which they step. For instance, in SOpenReturnSuccess, the function fresh_fd: history -> file_descr is used to generate a fresh unique file descriptor based on the previously generated file descriptors listed in the history h. Meanwhile, in SOpenReturnFail, the event EvOpen records with the error err that opening the file was unsuccessful.
The reader should observe that this operational semantics does not actually model any of the files being accessed as some form of state. Instead, the semantics models all the possible IO events that might be produced if actual files were to be accessed by programs written in $\lambda_{io}$.
To compute the reflexive-transitive closure of the step relation, we want to consider traces of events corresponding to the events produced by the sequence of steps. To maintain the well-formedness of events with respect to a particular history for each of the individual reduction steps, we define a notion of local traces, which are traces that are well formed with respect to a particular history from which they start. The well-formedness of a trace is calculated recursively.
let rec well_formed_local_trace (h:history) (lt:trace) =
match lt with | [] -> True
| EvOpen _ (Inl fd) :: tl -> fd == (fresh_fd h) /\ well_formed_local_trace (ev::h) tl
| ev :: tl -> well_formed_local_trace (ev::h) tl
This definition of well-formedness ensures that all the events in the local trace are well formed with respect to the given history and the events preceding them in the local trace. This refinement on traces greatly simplifies formalization and reasoning about the validity of a local trace. Monotonically increasing local traces are trivially well-formed, and the order of sequential traces is clear from histories by which the local traces are indexed. For example, a local trace lt2:local_trace (h++lt1) can only be appended to a local trace typed as lt1:local_trace h, giving lt1@lt2 : local_trace h. Here the append function (++) : (h:history) -> local_trace h -> history reverses the given local trace and prepends it to the history. Therefore, the most recent events appear at the beginning of the history and are in reverse order. For local traces, we use above the built-in append function (@) : #h:history -> lt:local_trace h -> lt’:local_trace (h++lt) -> local_trace h with the additional refinement that the resulting local trace is well formed with respect to the history h.
We then represent the reflexive-transitive closure of step as an inductive type steps indexed by two closed $\lambda_{io}$ expressions, a history, and a local trace modelling the events produced by the sequence of steps and indexed by the history where these steps start. The definition is mostly standard, with local traces getting concatenated when the corresponding reduction steps are concatenated.
type steps : closed_exp -> closed_exp -> h:history -> local_trace h -> Type =
| SRefl : e:closed_exp -> h:history -> steps e e h []
| STrans : #e1:_->#e2:_->#e3:_->#h:_ ->#oev:option(event_h h)->#lt:local_trace(h++(as_lt oev))->
step e1 e2 h oev -> steps e2 e3 (h++(as_lt oev)) lt -> steps e1 e3 h ((as_lt oev) @ lt)
We define the type e_beh e e h lt to represent steps e e h lt where e is an irreducible expression.
4.3. Syntax and semantics of IO
We represent the IO computations of the shallowly embedded IO language in a standard way, using a free monad, whose functor part io a is defined inductively by the following two cases.
type io (a:Type u#a) : Type u#a =
| Return : a -> io a
| Call : o:io_ops -> args:io_args o -> (io_res o args -> io a) -> io a
Elements of io a are computation trees, which intuitively either represent returning a pure value (Return x), or performing an operation o and proceeding as the continuation k (Call o args k). For example, the openfile fn operation is represented in this monad as the following F expression:
let openfile (fnm:string) : io (resexn file_descr) = Call OOpen fnm Return
The return and bind operations of this monad are defined in a standard way for a free monad:
let io_return (#a:Type) (x:a) : io a = Return x
let rec io_bind #a #b (c:io a) (k:a -> io b) : io b = match c with
| Return x -> k x
| Call o args ok -> Call o args (fun i -> io_bind (ok i) k)
We give semantics to computations represented using io in the tradition of Dijkstra monads (Maillard et al., 2019), in terms of a monad morphism from the io monad to some other suitable monad, which then specifies the computational behaviour of the io computations. The monad we use for the semantics is a predicate transformer monad (mapping postconditions to preconditions) specialised to IO computations and the traces they produce. The functor part of this monad is given by the type hist.
type hist_post (h:history) a = lt:local_trace h -> r:a -> Type0
type hist0 a = h:history -> hist_post h a -> Type0
let hist_post_ord (#h:history) (p1 p2:hist_post h a) = forall lt r. p1 lt r ==> p2 lt r
let hist_wp_monotonic (wp:hist0 a) =
forall h (p1 p2:hist_post h a). (p1 hist_post_ord p2) ==> (wp h p1 ==> wp h p2)
type hist a = wp:(hist0 a){hist_wp_monotonic wp}
Elements of hist a are monotonic predicate transformers that map a postcondition (a predicate on the local trace and return value of a computation) to a precondition (a predicate on the history trace of the computation). That the postcondition is only over the local trace of the computation and not the history plus the local trace, as one might naively expect, makes this monad akin to update monads (Ahman and Uustalu, 2013), and promotes more local and modular reasoning. An analogous, but slightly simplified predicate transformer monad for IO was also considered by Maillard et al. (2019).
The unit and return operations for this monad are defined analogously to other predicate transformer monads in the Dijkstra monads tradition (Maillard et al., 2019). One just needs to take care about correctly extending the history and local trace for the continuation in the bind operation.
let hist_return #a (x:a) : hist a = fun _ p -> p [] x
let post_shift #a (h:history) (p:hist_post h a) (lt:local_trace h) : hist_post(h++lt)a = fun lt r -> p (lt @ lt’) r
let hist_post_bind #a #b (#h:history) (kw : a -> hist b) (p:hist_post h b) : Tot (hist_post h a) =
fun lt r -> kw r (h ++ lt) (post_shift h p lt)
let hist_bind #a #b (w : hist a) (kw : a -> hist b) : hist b = fun h p -> w h (hist_post_bind kw p)
The return of this monad is the predicate transformer which requires the postcondition to hold for the empty local trace and the given pure value x. The bind composes predicate transformers. Due to the postconditions referring to local traces, we have to apply the postcondition p supplied to the composed predicate transformer to the concatenation of the local traces (see post_shift).
We consider two elements of hist a to be equal when they are pointwise equivalent.
let hist_equiv w1 w2 = forall h p. w1 h p <==> w2 h p
In order to give a semantics to IO computations, we define a monad morphism theta from computations represented in io a to predicate transformers in hist a, again in the tradition of Dijkstra monads (Maillard et al., 2019). Intuitively, the monad morphism theta returns a predicate transformer that computes for any postcondition the weakest precondition for the given computation. We define it as follows:
let op_wp (o:io_ops) (args:io_args o) : hist (io_res o args) =
fun h p -> True /\ (forall lt r. (io_post h o args r /\ lt == [op_to_ev o arg r]) ==> p lt r)
let rec theta #a (m:io a) : hist a = match m with
| Return x -> hist_return x
| Call o args k -> hist_bind (op_wp o args) (fun r -> theta (k r))
As we see, theta maps pure computations returning a value to the predicate transformer that requires the postcondition to hold for this value as the computed precondition. The operation calls are mapped to the composition (using hist_bind) of a predicate transformer corresponding to the operation (using op_wp, which returns a predicate transformer based on the pre- and postcondition of an operation) and the predicate transformer recursively computed from its continuation.
Finally, we use theta to specify the semantics of IO computations in terms of predicates on the local traces the computations produce and the values they return. Formally, the predicate fs_beh fs_e h we define is the strongest postcondition corresponding to the predicate transformer theta fs_e(Maillard et al., 2019)
let fs_beh #a (fs_e:io a) (h:history) : hist_post h a = fun lt res -> forall p. (theta fs_e) h p ==> p lt res
4.4. Quoted types at which programs are related
The next predicate characterizes the types we want to be able to extract, and, thus, the types of $\lambda_{io}$ and IO expressions that we want to be able to relate with the logical relations. It includes base types (unit, bool, and file_descr), function types (both pure and effectful), and product and sum types.
noeq type type_rep : Type -> Type =
| QUnit \$\ \$ : type_rep unit | QBool : type_rep bool | QFileDescriptor : type_rep file_descr
| QArr \$\ \ \ \$ : #a:Type -> #b:Type -> type_rep a -> type_rep b -> type_rep (a -> b)
| QArrIO : #a:Type -> #b:Type -> type_rep a -> type_rep b -> type_rep (a -> io b)
| QPair \$\ \ \$ : #a:Type -> #b:Type -> type_rep a -> type_rep b -> type_rep (a & b)
| QSum \$\ \$ : #a:Type -> #b:Type -> type_rep a -> type_rep b -> type_rep (either a b)
We define the type qType as a dependent pair between an F type and a proof that it satisfies the predicate, as t:Type0 & type_rep t. Formally, the environments and the typing relation are defined in terms of qType, but we hide it in the paper for readability.
4.5. Target-to-source and source-to-target logical relations for relating programs
To verify SEIO, we have to formally relate $\lambda_{io}$ programs with IO programs. At a high level, two programs are related if they have the same behavior—i.e., they evaluate to the same results and produce the same local traces. We relate a $\lambda_{io}$ program to an IO program using two logical relations, one for each direction. In the target-to-source relation, for all $\lambda_{io}$ behavior, there must exist corresponding IO behavior. In the source-to-target relation, we have the dual requirement.
As is standard (Ahmed, 2004), each logical relation is defined from a value relation and, because IO is a FGCBV language, two expression relations: a pure expression relation and an IO expression relation. The way that we define IO expressions, as either F values or F computations in the io monad, allows us to differentiate between pure and effectful expressions.
4.5.1. Logical relations for closed $\lambda_{io}$ and IO expressions
For the target-to-source logical relation, we define the value relation qt (h, fs_v, v) between histories h, IO values fs_v, and $\lambda_{io}$ values v by induction on (quoted) types in the natural way, as shown below. The only subtlety in this definition is that in the case of the two kinds of function types, the definition is given in the style of Kripke’s possible worlds semantics, where the function arguments are quantified and the applications considered in histories extended with arbitrary compatible local traces.
let (∋) (qt:qType) (h:history) (fs_v:qt._1) (v:value) = match qt._2 with
| QUnit -> fs_v == () /\ e == EUnit | QFileDescriptor -> e == EFileDescr fs_v | QString -> e == EString fs_v
| QBool -> (fs_v == true /\ e == Etrue) \/ (fs_v == false /\ e == Efalse)
| QArr qt1 qt2 -> let ELam e = e in forall (v:value) (fs_v:qt1._1) (lt_v:local_trace h).
qt1 (h++lt_v, fs_v, v) ==> (qt2 (h++lt_v, fs_f fs_v, subst_beta v e’))
| QArrIO qt1 qt2 -> let ELam e = e in forall (v:value) (fs_v:qt1._1) (lt_v:local_trace h).
qt1 (h++lt_v, fs_v, v) ==> (qt2 (h++lt_v, fs_f fs_v, subst_beta v e’))
(* QPair and QSum are defined as expected *)
The two expression relations are defined below. They specify that the behaviour of the $\lambda_{io}$ expression must imply a corresponding behaviour for the IO expression. The difference is that the pure relation requires the local trace to be empty. These relations express that any behaviour possible for a $\lambda_{io}$ expression that an IO expression is extracted to was already possible for the IO expression.
and (⊇) (qt:qType) (h:history) (fs_e:qt._1) (e:closed_exp) =
forall (e’:closed_exp) (lt:local_trace h). e_beh e e h lt ==> (t (h, fs_e, e’) /\ lt == [])
and (⫄) (qt:qType) (h:history) (fs_e:io qt._1) (e:closed_exp) =
forall (e’:closed_exp) (lt:local_trace h). e_beh e e h lt ==> (exists (fs_r:qt._1). t (h++lt, fs_r, e’) /\ fs_beh fs_e h lt fs_r)
For proving our security criterion in § 5, it is useful to restrict these relations to hold for all histories. For example, we define fs_e valid_contains e to mean that forall (h:history). t (h, fs_e, e). Analogously, we define fs_evalid_superset_val e and fs_e valid_superset_prod e to mean that forall (h:history). t (h, fs_e, e) and forall (h:history). t (h,fs_e, e).
For the source-to-target logical relation, the value relation qt ∈(h, fs_v, v) is defined analogously by induction on types, with the only difference between the two logical relations being how the pure and IO expression relations are defined that the cases for the two function types refer to.
let (∈) (qt:qType) (h:history) (fs_v:qt._1) (v:value) =
and (⊆) (t:qType) (h:history) (fs_e:qt._1) (e:closed_exp) =
exists (e’:closed_exp). e_beh e e h [] /\ t (h, fs_e, e’)
and (⫃) (qt:qType) (h:history) (fs_e:io qt._1) (e:closed_exp) =
forall (fs_r:qt._1) (lt:local_trace h). fs_beh fs_e h lt fs_r ==> (exists (e’:closed_exp). t (h++lt, fs_r, e’) /\ e_beh e e h lt)
The definitions are dual to the target-to-source logical relation defined above. Here we require that the behaviour of the IO expression must imply a corresponding behaviour for the $\lambda_{io}$ expression.
Similarly to the target-to-source relation, we also find it useful to restrict these relations to hold for all histories. For instance, we define fs_e valid_member_of e to mean that forall (h:history). t (h, fs_e, e).
4.5.2. Logical relations for open $\lambda_{io}$ and IO expressions
While closed expressions are evaluated at runtime, static type-checking rules operate on open expressions that might contain free variables. To use the logical relations we defined above on open expressions, we define type environments (mapping from variables to optional qTypes), IO evaluation environments (mapping from variables to IO values), and $\lambda_{io}$ evaluation environments (mapping from variables to $\lambda_{io}$ values—in other words, these are (parallel) substitutions for $\lambda_{io}$).
The type environment maps to optional qTypes because we use de Bruijn indices to represent variables. Therefore, every (natural number) index must have a corresponding entry, but only the free variables we seek to instantiate map to the appropriate qTypes. Both evaluation environments are indexed by the type environment. The $\lambda_{io}$ evaluation environment is additionally indexed by a boolean flag which specifies whether the substitution is in fact just a variable renaming.
We say that IO and $\lambda_{io}$ evaluation environments are logically related at a particular type environment tyG if the two evaluation environments map variables in tyG to related values.
let (∽) (#tyG:typ_env) (h:history) (fsGs:eval_env tyG) #b (eGs:gsub tyG b) =
forall (x:var). Some? (tyG x) ==> Some?.v (tyG x) (h, fsGs x, eGs x)
let (≍) (#tyG:typ_env) (h:history) (fsGs:eval_env tyG) #b (eGs:gsub tyG b) =
forall (x:var). Some? (tyG x) ==> Some?.v (tyG x) (h, fsGs x, eGs x)
Now, we can relate open $\lambda_{io}$ and IO expressions at a given type environment by specifying that for every related $\lambda_{io}$ and IO evaluation environment indexed by the type environment, the corresponding closed $\lambda_{io}$ and IO expressions must be related at the expected type. We present the definition for the target-to-source logical relation below. The source-to-target relation is analogous.
let (⊐) (#tyG:typ_env) (qt:qType) (fs_oe:eval_env tyG -> qt._1) (oe:exp) =
fv_in_env tyG oe /\
forall b (eGs:gsub tyG b) (fsGs:eval_env tyG) (h:history). fsGs ‘(∽) h eGs ==> t (h, fs_oe fsG, gsubst eGs oe)
The predicate fv_in_env specifies that all free variables in the $\lambda_{io}$ expression are mapped to Some qt by the type environment tyG. Open IO values are modelled as functions from an evaluation environment that closes them to the corresponding closed IO value. In the definition, fs_oe fsGs is the result of replacing all the free variables in fs_oe with their values under the IO evaluation environment fsGs, with gsubst eGs oe doing the same but for oe using the $\lambda_{io}$ evaluation environment eGs. For open $\lambda_{io}$ expressions and IO computations, their closed variants must be in the IO expression relation.
4.6. Compatibility of the logical relations with typing rules
As part of proving our security criterion in § 5.2, we prove in § 5.1 the fundamental property of logical relation for both of our logical relations, relating a source program to the corresponding compiled code. As is typical, our proofs of the fundamental property also proceed by induction on typing derivations, and to keep the proof manageable, one proves a number of compatibility lemmas showing that the logical relations satisfy rules analogous to the typing rules of the language.
Since our logical relations relate both pure and IO expressions, we need to prove three kinds of compatibility lemmas for each typing rule:
  1. (1)
    when the subexpressions are pure expressions and the composite expression is pure;
  2. (2)
    when some of the subexpressions are pure expressions and some are IO expressions and the composite expression is an IO expression;
  3. (3)
    when the subexpressions are IO expressions and the composite is also an IO expression.
We illustrate this with the three kinds of compatibility lemmas for the source-to-target logical relation corresponding to the typing rule for function application. The c1 lemma considers a pure arrow and a pure argument. The c2 lemma considers an arrow that takes a pure argument and returns an IO computation and a pure argument. The c3 lemma considers an IO arrow and an IO argument.
let c1 #tyG (#a #b:qType) (fs_f:eval_env tyG -> (a._1 -> b._1)) (fs_x:eval_env tyG -> a._1) (f x:exp)
: Lemma (requires fs_f f /\ fs_x x)
(ensures (fun fsG -> (fs_f fsG) (fs_x fsG)) EApp f x)
let c2 #tyG (#a #b:qType) (fs_f:eval_env tyG -> (a._1 -> io b._1)) (fs_x:eval_env tyG -> a._1) (f x:exp)
: Lemma (requires fs_f f /\ fs_x x)
(ensures (fun fsG -> (fs_f fsG) (fs_x fsG)) EApp f x)
let c3 #tyG (#a #b:qType) (fs_f:eval_env tyG -> io (a._1 -> io b._1)) (fs_x:eval_env tyG -> io a._1) (f x:exp)
: Lemma (requires fs_f f /\ fs_x x)
(ensures (fun fsG -> io_bind (fs_f fsG) (fun f -> io_bind (fs_x fsG) (fun x -> f x’))) EApp f x)
We sketch the proof of c3. Full details of the proofs are available in our artifact.
  1. (1)
    We consider arbitrary irreducible closed $\lambda_{io}$ expression e and local trace lt:local_trace h such that e_beh e e h lt, and we need to prove exists (fs_r:b._1). b (h++lt, fs_r, e’) /\ fs_beh fs_e h lt fs_r.
  2. (2)
    We then destruct the steps from e to e into the intermediate steps of the subexpressions.
    1. (a)
      For the lambda expression, we get e_beh f (ELam f’) h lt1 and e_beh e (EApp (ELam f’) x) h lt1
    2. (b)
      Since the reduction of the lambda expression corresponds to lt1:local_trace h, we reduce the argument expression starting from the history h++lt1. For the argument expression, we get e_beh x x (h++lt1) lt2 and e_beh (EApp (ELam f’) x) (subst_beta x f’) (h++lt1) lt2.
    3. (c)
      Finally, we get that e_beh (subst_beta x f’) e ((h++lt1)++lt2) lt3 and lt == lt1@lt2@lt3.
  3. (3)
    Next, we use the fact that relatedness of the $\lambda_{io}$ and IO evaluation environments is closed under history extension to apply the induction hypothesis and get information about the IO behavior corresponding to these three sequences of steps.
  4. (4)
    By unfolding the IO arrow, since we know that e_beh (subst_beta x f’) e ((h++lt1)++lt2) lt3 and lt == lt1@lt2@lt3, we get that exists (fs_r:b._1). b (h++lt, fs_r, e’) /\ fs_beh (fs_r_f fs_r_x) ((h++lt1)++lt2) lt3 fs_r.
  5. (5)
    Since we know that fs_beh fs_x (h++lt1) lt2 fs_r_x and
    fs_beh ((fun x -> fs_r_f x’) fs_r_x) ((h++lt1)++lt2) lt3, we use the fact that fs_beh is a monad morphism to get that fs_beh (io_bind (fs_x (fun x -> fs_r_f x’)) (lt2@lt3) fs_r).
  6. (6)
    Finally, we utilize this lemma once more with fs_beh fs_f h lt1 fs_r_f to get that
    fs_beh (io_bind fs_f (fun f -> io_bind fs_x (fun x -> f x’))) h (lt1@lt2@lt3) fs_r, which is exactly the desired fs_beh behavior. ∎
Other compatibility lemmas for the target-to-source relation always follow this general pattern:
  1. (1)
    Destruct e_beh e e h lt into subexpressions’ e_beh behavior.
  2. (2)
    Apply the induction hypothesis to get subexpressions’ fs_beh behavior.
  3. (3)
    Combine fs_beh behaviors to get the composite fs_beh behavior corresponding to e_beh e e h lt.
The compatibility lemmas for the source-to-target logical relation use the same overall idea as the target-to-source compatibility lemmas above, but they follow a dual proof pattern:
  1. (1)
    Destruct fs_beh fs_e h lt fs_r behaviour into subcomputations split at the io_binds in fs_e, so fs_beh m h lt1 fs_r_m and fs_beh (fun m -> k m’) (h++lt1) lt2 fs_r_k, where lt == lt1@lt2.
  2. (2)
    Apply the induction hypothesis to get subcomputations’ e_beh behavior.
  3. (3)
    Combine e_beh behaviors to get the composite e_beh behavior for e_beh fs_e h lt fs_r.
It is worth noting that destructing fs_beh behaviours is more complicated than destructing e_beh behaviours. While e_beh behaviours are defined inductively and can be effectively pattern-matched on, the fs_beh behaviours are defined in terms of logical implications (§ 4.3), leading to indirect reasoning about destructing local traces by instantiating them with suitably clever postconditions.
5. SEIO: Formally Secure Extraction
SEIO involves two steps: relational quotation and syntax generation. Relational quotation is a translation validation technique that involves a metaprogram, while syntax generation is a pure F function that we verify once and for all. In this section, we explain how we instantiate the compilation model of Abate et al. (2019) with SEIO (§ 5.1) to prove that it satisfies RrHP (§ 5.2).
5.1. Overview of the compilation model
When we instantiate the compilation model of Abate et al. (2019) with SEIO we cannot talk about the relational quotation metaprogram, since in F we cannot state any extrinsic property about a metaprogram. So we do this instantiation such that the final criterion we prove about the model directly relates the shallowly embedded program with its extraction produced by SEIO.
The source language is IO. A partial source program (denoted by $P$, of type progS) is a dependent pair containing (1) an F function in the io monad taking a context as argument and returning a Boolean, and (2) a typing derivation that proves the program is in the IO language. The partial source program and the source context share an interface (of type interface) that contains only the type of the context ct. A source context can be any F code with a type that can be quoted (§ 4.4). Source linking (linkS, denoted by $C^{S}[P]$) is function application.
noeq type interface = { ct : Type; qtyp: type_rep ct; }
(** Source **)
type progS (i:interface) = ps:(i.ct -> io bool) & (typing_oval empty (i.ct -> io bool) ps)
type ctxS (i:interface) = i.ct
type wholeS = io bool
let linkS (#i:interface) (ps:progS i) (cs:ctxS i) : wholeS = (dfst ps) cs
While each partial source program includes a typing derivation, it is there only to be able to do compilation (one can think of it as a compilation hint), but it is not relevant for the semantics of the partial program. One can see above that during source linking, the typing derivation is discarded, and whole source programs are just io computations. This is key to stating a secure compilation criterion that directly relates the shallowly embedded source program with its extraction.
(** Target **)
type progT (i:interface) = value
[…]
type wholeT = closed_exp
let linkT (#i:interface) (pt:progT i) (ct:ctxT i) : wholeT = EApp pt (dfst e)
Compilation is our syntax generation function, taking a source program $P$ and producing a target program denoted by $P\hskip-3.50006pt\downarrow$, which is a value in the target language $\lambda_{io}$ (of type progT). A target program can be linked securely with any syntactically typed target context (denoted by $C^{T}$, of type ctxT). To type target contexts, we introduce a simple syntactic typing relation, typingλ, defined as expected. For simplicity, when defining the typing relation for $\lambda_{io}$, we reuse the notion of types (§ 4.4) and typing environments from IO. We also have the same source and target interfaces. Finally, it is worth noting that the extracted program $P\hskip-3.50006pt\downarrow$ is not necessarily syntactically well typed, but we prove it is semantically well typed as part of the logical relation proofs below. The target of our extraction is basically untyped, similarly to Malfunction (Forster et al., 2024), as opposed to the default F and Rocq extraction that has to work hard to work around the limitations of a syntactic type system (Letouzey, 2004).
Compilation is defined recursively on the typing relation of IO and generates $\lambda_{io}$ syntax. We prove in F that the source and compiled programs are related by both logical relations. Since compilation returns a $\lambda_{io}$ value, we show that the source and compiled program are related for all histories by the two value relations (by valid_contains and valid_member_of). We note that these results are the appropriate notions of the fundamental property of logical relations for our two logical relations. Thus, as is common, their proofs involve applying the relevant compatibility lemmas discussed in § 4.6.
Finally, we give semantics to whole source and target programs using $\leadsto_{S}$ and $\leadsto_{T}$, which are defined in terms of fs_beh (§ 4.3) and e_beh (§ 4.2) starting from the empty history ([]).
[…]
5.2. Robust Relational Hyperproperty Preservation (RrHP)
We show that SEIO satisfies Robust Relational Hyperproperty Preservation (RrHP), the strongest secure compilation criterion of Abate et al. (2019). RrHP not only implies full abstraction (i.e., preserving observational equivalence), but also ensures the preservation of trace properties and hyperproperties (like noninterference) against adversarial contexts.
Theorem 5.1 (Robust Relational Hyperproperty Preservation (RrHP)).
$\forall I^{S}.\ \forall C^{T}.\ \exists C^{S}.\ \forall P:\mathsf{prog}^{S}\ I^{S}.\ \forall t.\ (C^{T}[P\hskip-3.50006pt\downarrow]\leadsto_{T}t\Leftrightarrow C^{S}[P]\leadsto_{S}t)$
The statement of RrHP uses an equivalence, so we prove each direction separately. In both cases, to instantiate the existential for source contexts, we use a back-translation function (denoted by $C^{T}\hskip-3.50006pt\uparrow$) from target contexts to source contexts. Since target contexts are syntactically well-typed, back-translation is defined recursively on the typing relation of $\lambda_{io}$. We prove in F that the source context obtained by back-translation is related to the original target context with respect to both our logical relations. These results are another instance of the fundamental property of logical relations for our two relations, so the proofs again rely on compatibility lemmas discussed in § 4.6.
To prove the right-to-left direction of RrHP ($\Rightarrow$), we use the target-to-source logical relation that relates each target behavior to a source behavior (§ 4.5).
Theorem 5.2 (RrHP $\Rightarrow$).
$\forall I^{S}.\ \forall C^{T}.\ \exists C^{S}.\ \forall P.\ \forall t.\ C^{T}[P\hskip-3.50006pt\downarrow]\leadsto_{T}t\Rightarrow C^{S}[P]\leadsto_{S}t$
Proof sketch.
We back-translate the target context to get a source context to instantiate the existential. Since compilation is verified and $P$ is a value, we know that $P$ valid_contains $P\hskip-3.50006pt\downarrow$. Since back-translation is verified and $C^{T}$ is a value, we know that $C^{T}\hskip-3.50006pt\uparrow$ valid_contains $C^{T}$. Since our programs are functions, it follows that $(P~C^{T}\hskip-3.50006pt\uparrow)$ valid_superset_prod (EApp $P\hskip-3.50006pt\downarrow C^{T}$). Since linking is function application, this is all we need. ∎
This direction of RrHP already implies Robust Trace Property Preservation (RrTP), which is a weaker criterion that RrHP, but still one of the strongest secure compilation criteria of Abate et al. (2019). They show that RrTP implies full abstraction, and also ensures the preservation of trace properties and subset-closed hyperproperties against adversarial contexts.
To prove the other direction of RrHP, we use the source-to-target logical relation (§ 4.5).
Theorem 5.3 (RrHP $\Leftarrow$).
$\forall I^{S}.\ \forall C^{T}.\ \exists C^{S}.\ \forall P.\ \forall t.\ C^{S}[P]\leadsto_{S}t\Rightarrow C^{T}[P\hskip-3.50006pt\downarrow]\leadsto_{T}t$
Proof sketch.
This is dual to the proof of the Theorem 5.2, just this time using the logical relation that relates each source behavior to a target behavior. ∎
6. SEIO in action
This section provides more details on how to run the running example. In § 6.1, we first present a concrete validate function and agent definitions that make this validate hold or not. Later, in § 6.2, we briefly explain how to connect our framework with other tools to obtain an executable file that allows us to test SEIO in action.
6.1. A verified wrapper for unverified AI agents
We tested our framework with the verified wrapper for agents we presented in § 2. For the validate construction we used a very simple condition:
let validate olds task news = eq_string task news
that tests whether the task description coincides with the new content of the file. The function eq_string is assumed as a primitive constant of $\lambda_{io}$. Other string operations, such as concatenation, could be added and treated in a similar fashion.
The verified wrapper can be run together with agents written directly as $\lambda_{io}$ expressions. Even when the primitives present currently in $\lambda_{io}$ are limited, we still tested a number of things. These are some agents we have tried that work as intended:
(* bad: this agent does nothing *)
let lazy_agent : exp = ELam (ELam EUnit)
(* good: this agent writes the expected string on the file *)
let write_agent : exp =
ELam (ELam (ECase (EOpen (EVar 1)) (EApp (ELam (EClose (EVar 1))) (EWrite (EVar 0) (EVar 1))) EUnit ))
(* bad: this agent writes the string to the file twice *)
let write_twice_agent : exp =
ELam (ELam (ECase (EOpen (EVar 1))
(EApp (ELam (EApp (ELam (EClose (EVar 2)))
(EWrite (EVar 1) (EVar 2))))
(EWrite (EVar 0) (EVar 1))) EUnit))
6.2. Unverified compilation step from $\lambda_{io}$ to $\lambda_{\square}$
[…]
Figure 4. Compilation pipeline from $\lambda_{io}$ to executable.
To build an actual executable from the programs we can write in SEIO, we complement the proposed extraction mechanism with an unverified step compiling from $\lambda_{io}$ to $\lambda_{\square}$ (lambdabox), and then using the tooling provided by the Peregrine project (Chapman et al., ) to compile to OCaml (Malfunction (Forster et al., 2024; Dolan, 2016)) to obtain code that can be later linked with a runtime in OCaml.
The $\lambda_{\square}$ language was introduced in the context of program extraction for Rocq (Letouzey, 2004). It is a variant of untyped $\lambda$-calculus with a construction $\square$ (box) that replaces proofs and types that are computationally irrelevant. In our use case, we do not erase proofs, as $\lambda_{io}$ already does not contain any, and we use it just as an intermediate language that allows us to use the rest of the compilation pipeline provided by Peregrine.
Using Peregrine, the $\lambda_{\square}$ code is translated to Malfunction, which is a low-level untyped program representation that can be used to compile to OCaml objects. The Peregrine framework uses this program representation as its output for OCaml. This output can later be compiled using Malfunction tooling (Dolan, 2016) to generate OCaml objects, which can be linked with other OCaml objects that implement IO primitives such as openfile, readfile, etc. as well as other primitives of $\lambda_{\square}$ such as string equality. This compilation and linking sequence is presented in Figure 4.
7. Related work
Certifying extraction
Work on CakeML synthesis (Myreen and Owens, 2014; Abrahamsson et al., 2020) shows how starting from a shallowly embedded program in HOL one can synthesize an equivalent CakeML (Kumar et al., 2014) program (a subset of Standard ML). Their synthesis happens outside of HOL, in Standard ML, and it produces a deeply embedded program in HOL and a correctness proof that has to be accepted by HOL, which is a form of translation valuation. They show how to synthesize CakeML programs from HOL functions that feature polymorphism, first-order references, arrays, exceptions, and IO operations. We think we can extend SEIO with these features by following their lead to obtain certifying extraction from F with a smaller TCB and stronger security guarantees than in their work (Myreen and Owens, 2014; Abrahamsson et al., 2020). Another difference comes from the languages we use, they work in HOL, a simply typed language, while we work in F, a dependently typed language, which enabled us to factor out relational quotation as a separate step.
Œuf (Mullen et al., 2018) is a compiler with end-to-end guarantees from a simply typed subset of Gallina (the language of Rocq) to Assembly code. They use translation validation for quotation, which produces a deeply embedded program. They validate this by showing in Rocq that the denotation of the deep embedding is equal to the Gallina program. They define denotation as a pure Rocq function that takes a deeply embedded program and produces a shallowly embedded one. This is similar to what happens in relational quotation, where the typing derivation also constructs a program that is checked to be equal with the program that is quoted. From our experience, using a typing relation works better than separately defining a deep embedding and a denotation function. For example, it is more natural to keep in the typing derivation extra information necessary to reconstruct the shallow embedding, information which may have no place in a standard deep embedding. We used relational quotation for a language with the IO effect, while Œuf has no support for effects.
Forster and Kunze (2019) present a certifying extraction from pure Rocq functions of simple polymorphic types to an untyped call-by-value $\lambda$-calculus. They present a plugin for Rocq (a metaprogram) that given a shallowly embedded program, produces a deeply embedded program. The plugin comes with a series of tactics that automate the process of validation that the two programs are related. They have no support for effects, while SEIO supports IO.
Pit-Claudel et al. (2022) present a Rocq framework that treats compilation as a proof-search problem—i.e., finding a deeply embedded low-level program that is in a relation with the shallowly embedded program. This is also a form of translation validation, since the search is done using tactics (metaprograms) whose result has to be type-checked by Rocq. The framework may not know how to automatically compile a program, in which case it shows where it got stuck, and the proof engineer can take control. They illustrate the versatility of the framework by showing how to compile Gallina programs using terminating loops, various flat data structures such as mutable first-order cells and arrays, nondeterminism, and IO. Pit-Claudel et al. (2022) present their framework as translation validation tool for specific performance-critical programs, directly targetting a low-level program. We on the other hand try to minimize the reliance on translation validation.
The work above and our relational quotation share a challenge that comes from trying to relate a shallowly embedded program with a deeply embedded program: one is able to relate only what is expressible by a user defined relation (e.g., pattern matching is hard to do: when synthesizing CakeML (Abrahamsson et al., 2020; Myreen and Owens, 2014), they first automatically generate the relation based on the datatypes used). We came up with the typing relation for shallow embeddings after reading work on certifying extraction (Myreen and Owens, 2014; Abrahamsson et al., 2020; Pit-Claudel et al., 2022), but, to our knowledge, it does not exist per se in literature. In terms of similar looking relations, we are aware only of work on canonical structures (Gonthier et al., 2013).
Logical relations
Our logical relations, and their related proofs, follow the logical relation for semantic typing presented by Ahmed (2004) (§2.2) for a $\lambda$-calculus extended with dynamically allocated immutable cells. We differ from it in three ways: our logical relations are cross-language (Neis et al., 2015; Ahmed, 2015) and asymmetric, connecting a shallow and a deep embedding. They relate a program of a FGCBV language to one of a CBV language, and relates two trace-producing semantics.
Asymmetric relations are common in work on certified extraction, and Forster and Kunze (2019) and work on CakeML synthesis (Abrahamsson et al., 2020; Myreen and Owens, 2014) also use logical relations. The most complex relation is used in CakeML synthesis: it supports multiple effects (first-order state, exceptions and IO), polymorphic functions and total recursive functions, and relates evaluation semantics. Our logical relation is different because it relates two trace-producing semantics and it implies the strongest secure compilation criterion (Abate et al., 2019) (RrHP).
Logical relations were used to prove fully abstract compilation (New et al., 2016; Devriese et al., 2016; Patrignani et al., 2019) and Abate et al. (2019) later show that they can also be used to prove their stronger RrHP secure compilation criterion for a simple translation from a statically typed language to a dynamically typed one.
Verified extraction
CakeML (Kumar et al., 2014) is a subset of Standard ML that has a verified compiler that can target ARMv6, ARMv8, x86-x64, MIPS-64, RISC-V and Silver RSA architectures. Together with their synthesis from HOL to CakeML, they have end-to-end guarantees of correctness from the original program written in HOL to the low-level code that targets a specific architecture. It would be interesting to extend SEIO to target CakeML with a verified secure compilation step.
Hupel and Nipkow (2018) and Nezamabadi et al. (2026) present verified extraction from Isabelle/HOL and Dafny to CakeML, but there is no formal connection between the formalized meta-theory of Isabelle/HOL or Dafny and the languages themselves.
Forster et al. (2024) present a verified compiler from Gallina to Malfunction (an intermediate untyped language of OCaml). Their compiler is based on the existing MetaRocq (Sozeau et al., 2025) project that formalizes Rocq in itself. They prove a realizability relation that guarantees that the extracted untyped code operationally behaves as the initial Rocq program. They do not support effectful code, and quotation is neither verified nor certifying.
CertiCoq (Anand et al., 2017) and ConCert (Annenkov et al., 2022; Meier et al., 2025) are two partially verified compilers for Rocq that target C, Elm, Rust, WebAssembly and blockchain languages. They also rely on unverified quotation.
Secure compilation of verified code
There are two secure compilation frameworks for F, SCIO (Andrici et al., 2024) for verified IO programs, and SecRef (Andrici et al., 2025) for verified stateful programs. They are both proved to satisfy RrHP, but they are not fully verified: they rely on the unverified extraction mechanism of F to OCaml. SCIO and SecRef look at how to convert refinement types, pre- and post-conditions into dynamic checks using higher-order contracts and a reference monitor. SEIO is orthogonal to them, since it focuses on providing guarantees for the extraction step itself. Beyond such secure extraction from proof assistants, there are many other applications of secure compilation (Patrignani et al., 2019; Abate et al., 2019).
8. Conclusion and future work
We have shown that formally secure extraction of shallowly embedded effectful programs can be achieved by using translation validation only for a first step—relational quotation—while using a once-and-for-all verified syntax generation function for the rest. We have illustrated this general idea for F by developing SEIO, a framework that extracts F programs with IO while providing machine-checked secure compilation guarantees (RrHP).
In the future, we would like to connect SEIO with SCIO  (Andrici et al., 2024), the previous secure compilation framework for verified F programs with IO, already discussed at the end of § 7 above. This raises several challenges though: First, while SCIO turns some of the refinement types and pre-/post-conditions on the interface between program and context into dynamic checks, the compiled programs produced by SCIO can still make use of these features (e.g., internally). We believe relational quotation can be extended to these features, but this is made challenging by the fact that in F logical facts from the context (e.g., refinements on arguments or function pre-conditions) can be used automatically by the SMT solver, without any syntactic marker in the program. Second, SCIO works around F’s lack of general effect polymorphism by making use of a restricted form called flag polymorphism to model that the context is prevented from directly calling the IO primitives to avoid the access control done by a reference monitor. Third, SCIO makes use of F’s effect mechanism, and while effects in F can be reified to their underlying monadic representation (Grimm et al., 2018), support for reasoning about the obtained monadic computations is at the moment still too limited to be able to do the extrinsic proofs that would be needed to connect SCIO and SEIO.
It is worth noting that these challenges are specific to F and the general idea of using relational quotation to split extraction into two distinct phases could also be applied to more standard dependently typed proof assistants like Rocq and Lean. For instance, it would be interesting to combine and extend the ideas of both SCIO and SEIO and apply them to Loom (Gladshtein et al., 2026), a recent verification framework for effectful Lean programs, which like F is based on Dijkstra monads (Maillard et al., 2019). This would also require extending relational quotation to more effects, but prior work on CakeML synthesis (Myreen and Owens, 2014; Abrahamsson et al., 2020) and relational compilation (Pit-Claudel et al., 2022) suggest this should be possible.
Acknowledgements.
We are grateful to Guido Martínez, Yong Kiam Tan and Jérémy Thibault for the insightful discussions. This work was in part supported by the Sponsor European Research Council https://erc.europa.eu/ under ERC Starting Grant SECOMP (Grant #715753), by the German Federal Ministry of Education and Research BMBF (grant 16KISK038, project 6GEM), and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the Excellence Strategy of the German Federal and State Governments – EXC 2092 CASA - 390781972. Exequiel Rivas was supported by the Estonian Research Council starting grant PSG749. Danel Ahman was supported by the Estonian Research Council grant PRG2764.